A rope whose working strength is 2,000 N is used to tow a 1,000 kg car up a 10 ° incline. Find the maximum acceleration that can be given to the car. The coefficient of friction (µk) between the road and tires is approximately 0.6.
1
Expert's answer
2020-12-08T06:27:43-0500

Answer

Downward force due to weight

"F=mgsin\theta=1000\times 9.8\times sin10°=1666N"

kinetic friction force on car

"F_2=\mu mgcos\theta=1000\times 9.8\times cos10°=5762.4N"

Minimum force required to tow car in upward direction is

"F_{minimum}=F+F_1=5762+1666=7428.4N
\\T<F_{minimum} \\so\\acceleration\\a=0m/s^2"